OpenStudy (anonymous):

PLEASE. :) Find the 19th term for the arithmetic sequence. 1, -5, -11, -17, ... a19= ??

OpenStudy (anonymous):

I think u will again ask the answer.. a1=1, d=-6 and n=19 a19 = a1 + (n-1)d = 1 + 18(-6) = 1-108 = -107

OpenStudy (anonymous):

haha. u're good. thank u so much, abhimanyuPudi :))

OpenStudy (anonymous):

@AbhimanyuPudi

OpenStudy (anonymous):

welcome :-)